# Representative Ranking for Deliberation in the Public Sphere

### v1

@@ 3.4 The Price of Justified Representation
Justified representation need not be compatible with the objectives of general scoring rules. [^6] We show how JR conflicts with diverse approval in Fig. B.1. We formally analyze the question: what price must be paid in the total score if we require that selected items satisfy JR? We define the optimal JR set $S^{*}_{JR}(\mathcal{I}_{m,\mathcal{A}_{n},k},f)$ as the set that maximizes the score among all sets that satisfy justified representation [^7] Our results hold irrespective of random tie-breaking.:
@@ 5 The Price of JR in a Clustered Setting
We show that when $\gamma$ is small and the groups display high within-group homogeneity, the price of JR is low. We operationalize within-group homogeneity in two distinct ways, showing that the price of JR remains consistently low under both interpretations.
@@ 5.2 Homogeneity as Low Dispersion
Operationalizing homogeneity as cohesiveness was illustrative, but in real-world scenarios, it may be unlikely for all the members in a divided group to approve on one comment in common. For a more realistic model, we now operationalize within-group homogeneity in the context of a common statistical model of preferences. In particular, we assume that users’ preferences over items are drawn from a Mallows mixture model (Awasthi et al., 2014; Liu and Moitra, 2018), where each divided group has its own component in the mixture. We then operationalize homogeneity by the level of dispersion $\phi\in[0,1]$ of individual preferences from their group-specific central ranking. In this setting, we find that with high probability, the price of JR is at most $O\left(k/\left(k-\gamma\left\lceil(\log k)/\log(1/\phi)\right\rceil\right)\right)$, which approaches our previous bound of $O(k/(k-\gamma))$ as within-group homogeneity increases, i.e., as $\phi\rightarrow 0$. Let us begin by defining the Mallows model.
@@
A sample from the Mallows model represents a user $u$’s ranked preference $\pi_{u}$ over all items. For the approval voting setting, we assume that user approves of the top $\tau$ items in their ranking $\pi$, i.e., $A_{u}=\pi_{u}^{-1}([\tau])$. We can obtain an analytical upper bound on the price of JR.
@@
Figure 2: Representation when selecting $k=8$ Remesh comments, with and without a JR constraint. An individual who approves of at least one comment in a set $S$ is said to be represented. The JR constraint significantly increases the number of individuals who are represented. The prices of JR are reported with the shorthand $P(f)$ for $P^{\text{Greedy}}(\mathcal{I}_{m,\mathcal{A}_{n},8},f).$ Due to space constraints, the figure only shows the five questions for which $S^{*}$ did not satisfy JR under diverse approval. Fig. E.2 shows results for the remaining five questions.
@@ 6 Experiments
The GreedyCC algorithm and its price. Solving for the optimal JR set $S^{*}_{JR}$ is NP-hard in general (Bredereck et al., 2019a; Elkind et al., 2022). For small instances, we may use an ILP formulation, but for large instances (e.g. on social media platforms), we would expect practitioners to resort to approximation algorithms. To understand the effects we might expect in practice, we too employ an approximation algorithm in all our experiments, along with an associated approximate price of JR. Specifically, we use the GreedyCC algorithm (Elkind et al., 2022, 2023) to find a set $S^{\text{Greedy}}_{JR}$ that satisfies JR and approximately maximizes the score function $f$. In brief, our implementation of GreedyCC greedily adds items to the selected set until a $n/k$-justifying set is found; the remaining items are then selected to maximize the score function $f$. A full description of the algorithm is given in Alg. 16.
@@
For each question in the Remesh dataset [^13] We filter the data to remove any empty or duplicate comments. and each score function $f$, we compare the optimal set $S^{*}(f)$ of $k=8$ comments (without a JR constraint) and the set $S^{\text{Greedy}}_{JR}(f)$ which optimizes the score function while ensuring JR through the GreedyCC approximation algorithm (Alg. 16).
@@ 6.1 Results
Ranking with Perspective API always satisfied JR. Without explicitly enforcing JR, the engagement $S^{*}(f_{\text{eng}})$, diverse approval $S^{*}(f_{\text{DA}})$, and the Perspective API set $S^{*}(f_{C})$ satisfied JR 30%, 50%, and 100% of the time, respectively. This was surprising as the Perspective API does not explicitly aim to satisfy JR. Moreover, both diverse approval and the Perspective API classifiers are meant to“bridge” across groups (Ovadya and Thorburn, 2023a). However, Perspective API satisfied JR at a much higher rate and also provided better overall representation (Fig. 2). This could be because diverse approval focuses on bridging across a few pre-defined groups (in this case, the three political groups), whereas the bridging attributes targeted by the Perspective API might appeal to a broader range of groups. [^14] This might also result from the Remesh inferred votes that may predict higher probability for, e.g., longer texts which also get higher Perspective API scores.
@@
The relatively low prices may be because in all questions GreedyCC was able to find a $n/k$-justifying set with only at most two comments (Table E.3). A small $n/k$-justifying set would be expected in settings where user opinions clusters into only a few groups (the setting we analyze in Sec. 5.2).
@@ 7 Discussion
Limitations. In our work, we implemented a proof-of-concept that illustrates the benefits of incorporating representation constraints in conjunction with existing algorithmic interventions for comment ranking. However, there are still additional questions that should be considered before deploying it on a real-world platform.
First, the JR representation constraint depends upon an approval matrix that specifies which users approve of which comments. However, it is not always clear which users should be included in the approval matrix. For example, should it be all users on the platform or only those who saw the post? The answer to this question will likely be different for different platforms.
@@
Future work. In this work, we focused on JR because it is the foundational axiom that other extensions build upon. However, in future work, it would be interesting to explore strengthened versions of JR such as EJR (Aziz et al., 2017), EJR+ (Brill and Peters, 2023) or BJR (Fish et al., 2024). EJR+ seems like an especially promising direction since the problem of optimizing a linear score function subject to EJR+ can be solved through an ILP. Additionally, our work focused on the classical JR setting, where a set of $k$ top comments is selected. However, it could be interesting to explore extensions of JR that apply directly to the ranking setting (Skowron et al., 2017; Israel and Brill, 2024), although it is not immediately clear how to integrate a scoring function into these extensions.
@@ Acknowledgements
We thank Hoda Heidari, Seth Lazar, Aviv Ovadya, Ariel Proccacia, Luke Thorburn, and Glen Weyl for their feedback on this project.
@@ Appendix A How does JR compare to demographic-based representation?
It is worth discussing how our JR-based approach differs from other, perhaps more familiar, notions of demographic or social representation from the algorithmic fairness literature [Barocas et al., 2023, Chasalow and Levy, 2021]. JR is a bottom-up measure where the groups that are represented depend entirely on the users’ own approval of different comments. This has several advantages for the comment ranking setting.
First, is the feasibility for real-world applications. In algorithmic fairness, groups are typically pre-defined based on a demographic, such as race or religion. In contrast to JR, the requirement of (inference of) user demographic labels make applying many algorithmic fairness methods infeasible or difficult to implement in industry settings, including on social media platforms, because of conflicts with privacy and legal constraints [Holstein et al., 2019, Veale and Binns, 2017].
@@
Third, JR may accommodate intersectionality [Crenshaw, 1991] in a more natural way than many algorithmic fairness approaches. Algorithmic fairness approaches to intersectionality attempt to provide guarantees to a wide set of subgroups [Gohar and Cheng, 2023], e.g., by defining intersectional groups as the combination of different demographic attributes (e.g. age, gender, race) [Kearns et al., 2018]. However, this approach can lead to issues where the subgroups no longer correspond to meaningful entities. For instance, intersecting many dimensions can result in subgroups that are too specific and lack a meaningful reason for grouping (e.g., ”Jewish white males aged 18-35 who have a bachelor’s degree and live in a rural area”). In contrast, JR focuses on cohesive groups—groups of users who can agree on at least one comment in common—thereby inherently embedding some requirement for meaningfulness.
@@ Appendix B An Example of Conflict between Diverse Approval and JR
Figure B.1: When diverse approval fails at representativeness An approval profile with $n=12,m=5$ and $k=3.$ There are $\gamma=2$ cohesive groups $G_{1}=\{A,B,C,D,E,F\}$ and $G_{2}=\{G,H,I,J,K,L\}$ such that $|G_{1;2}|=6>n/k.$ Note that the sub-groups such as $\{A,B,C,D\}$ and $\{I,J,K,L\}$ are also cohesive and larger than $n/k$, so they ought to also be represented by JR. In turn, any JR set $S$ contains $\{1,2\}.$ On the other hand, the comments’ maximin diverse approval scores (across the groups $G_{1}$ and $G_{2}$) are $f(1)=f(2)=0$ and $f(3)=f(4)=f(5)=1/6.$ In turn, the set with the highest score is $S^{*}=\{3,4,5\}.$
@@ C.2 Proof of Theorem 1
We first prove the following lemma. This statement can be found in Bredereck et al. [2019b] (Theorem 3), but the authors do not provide a proof for the result in their paper so we write one below.
@@ C.4 Proof of Theorem 3
There are several generative interpretations of the Mallows model [Lu and Boutilier, 2014] but one that is particularly useful for analyzing homogeneity is based on the Repeated Insertion Model (RIM) [Doignon et al., 2004]. In our analysis, we make use of a modified version of RIM in which items from the reference ranking are inserted starting with the last ranked item and ending with the first ranked item. It can be shown this bottom up RIM sampling induces the same distribution as the usual top down RIM sampling.
@@ D.1 The GreedyCC approximation algorithm
First, we review the GreedyCC algorithm used in both sets of experiments. Given that solving for the optimal JR set $S^{*}_{JR}$ is NP-hard in general [Bredereck et al., 2019a, Elkind et al., 2022], we would expect practitioners to resort to approximation algorithms. Thus, to understand the effects we might expect in practice, in all our experiments, we too employ an approximation algorithm, along with an associated approximate price of JR. Specifically use the GreedyCC algorithm, which has also been used in prior work [Elkind et al., 2022], to find a set $S^{\text{Greedy}}_{JR}$ that satisfies JR and approximately maximizes the score function $f$. In brief, our implementation of GreedyCC greedily adds items to the selected set until a $n/k$-justifying set is found; the remaining items are then selected to maximize the score function $f$. A full description of the algorithm is outlined in Alg. 16.
@@ D.2 Mallows Mixture Model Simulations
$\mathcal{M}_{2}([\phi,\phi],[\sigma_{1},\sigma_{2}],[1/2,1/2])$ that simulates a polarized environment with two groups. We draw inspiration from Esteban and Ray [1994]’s axiomatic framework for measuring polarization. Esteban and Ray [1994] where they consider polarization to be present if three intuitive properties are satisfied. We operationalized these three properties in our simulation as follows:
@@ Results.
Second, we find that the price of GreedyCC is higher for diverse approval than for engagement. For diverse approval, some instances result in a price of GreedyCC that nearly our theoretical bound from Theorem 3. The gap in price between diverse approval and engagement is consistent with our result in Theorem 1, which found that the price of JR for maximin diverse approval $P(k,f_{\text{MDA}})$ equals $k$, whereas Elkind et al. [2022] demonstrated that the price for engagement $P(k,f_{\text{eng}})$ is $\Theta(\sqrt{k})$. These results were shown, in the general setting, without any restrictions on the approval profiles. On the other hand, the theoretical bound we derived for the polarized setting in Theorem 3, and plotted in Figure D.1, holds for all score functions. But given the empirical disparity in price between diverse approval and engagement, it seems plausible that a similar score-function-specific gap could be identified even in the polarized setting.
@@ E.2 Perspective API Classifiers
Figure E.1: Scores from the Perspective API
@@ E.4 Size of the $n/k$-justifying Set Found by GreedyCC
In Section 5.2, we show that when the user population clusters into a few groups, then the price of JR will be low because it is possible to find a small $n/k$-justifying set (and thus, the remaining items for the set can be chosen to maximize the score function). Indeed, this logic is how the GreedyCC approximation algorithm that we use operates. As detailed in Footnote 16, the GreedyCC proceeds in two stages. In the first stage, it finds an $n/k$-justifying set. In the second stage, it selects the remaining comments that maximize the score function. Thus, the smaller the $n/k$-justifying set, the lower the price of GreedyCC will tend to be.

### v2

@@ 3.4 The Price of Justified Representation
Justified representation is not guaranteed to be compatible with the objectives of general scoring rules. [^7] We show how JR conflicts with diverse approval in Fig. B.1. We formally analyze the question: what price must be paid in the total score if we require that selected items satisfy JR? We define the optimal JR set $S^{*}_{JR}(\mathcal{I}_{m,\mathcal{A}_{n},k},f)$ as the set that maximizes the score among all sets that satisfy justified representation [^8] Our results hold irrespective of random tie-breaking.:
@@ 5 The Price of JR in a Clustered Setting
We show that when $\gamma$ is small and the groups display high within-group homogeneity, the price of JR is low. We operationalize within-group homogeneity in two distinct ways, showing that the price of JR remains consistently low under both interpretations. These results are notable, because, in practice, when bridging interventions are implemented, (a) $\gamma$ is small and (b) divided groups are often determined by clustering or matrix factorization of the user approval matrix, resulting in groups that are relatively distinct and homogeneous (Wojcik et al., 2022; Small et al., 2021b).
@@ 5.2 Homogeneity as Low Dispersion
Operationalizing homogeneity as cohesiveness was illustrative, but in real-world scenarios, it may be unlikely for all the members in a divided group to approve on one comment in common. For a more realistic model, we now operationalize within-group homogeneity in the context of a common statistical model of preferences. In particular, we assume that users’ preferences over items are drawn from a Mallows mixture model (Awasthi et al., 2014; Liu & Moitra, 2018), where each divided group has its own component in the mixture. We then operationalize homogeneity by the level of dispersion $\phi\in[0,1]$ of individual preferences from their group-specific central ranking. In this setting, we find that with high probability, the price of JR is at most $O\left(k/\left(k-\gamma\left\lceil(\log k)/\log(1/\phi)\right\rceil\right)\right)$, which approaches our previous bound of $O(k/(k-\gamma))$ as within-group homogeneity increases, i.e., as $\phi\rightarrow 0$. Let us begin by defining the Mallows model.
@@
A sample from the Mallows model represents a user $u$’s ranked preference $\pi_{u}$ over all items. For the approval voting setting, we assume that user approves of the top $\tau$ items in their ranking $\pi$, i.e., $A_{u}=\pi_{u}^{-1}([\tau])$. We can obtain the following analytical upper bound on the price of JR, which is independent of any particular algorithmic strategy.
@@
Fig. 2: Representation when selecting $k=8$ Remesh comments, with and without a JR constraint. An individual who approves of at least one comment in a set $S$ is said to be represented. The JR constraint significantly increases the number of individuals who are represented. The prices of JR are reported with the shorthand $P(f)$ for $P^{\text{Greedy}}(\mathcal{I}_{m,\mathcal{A}_{n},8},f).$ Due to space constraints, the figure only shows the five questions for which $S^{*}$ did not satisfy JR under diverse approval. Fig. E.2 shows results for the remaining five questions.
@@ 6 Experiments
The GreedyCC algorithm and its price. Solving for the optimal JR set $S^{*}_{JR}$ is NP-hard in general (Bredereck et al., 2019; Elkind et al., 2022). For small instances, we may use an ILP formulation, but for large instances (e.g. on social media platforms), we would expect practitioners to resort to approximation algorithms. To understand the effects we might expect in practice, we too employ an approximation algorithm in all our experiments, along with an associated approximate price of JR. Specifically, we use the GreedyCC algorithm (Elkind et al., 2022, 2023) to find a set $S^{\text{Greedy}}_{JR}$ that satisfies JR and approximately maximizes the score function $f$. In brief, our implementation of GreedyCC greedily adds items to the selected set until a $n/k$-justifying set is found; the remaining items are then selected to maximize the score function $f$. A full description of the algorithm is given in Alg. 18.
@@
For each question in the Remesh dataset [^14] We filter the data to remove any empty or duplicate comments. and each score function $f$, we compare the optimal set $S^{*}(f)$ of $k=8$ comments (without a JR constraint) and the set $S^{\text{Greedy}}_{JR}(f)$ which optimizes the score function while ensuring JR through the GreedyCC approximation algorithm (Alg. 18).
@@ 6.1 Results
Ranking with Perspective API always satisfied JR. Without explicitly enforcing JR, the engagement $S^{*}(f_{\text{eng}})$, diverse approval $S^{*}(f_{\text{DA}})$, and the Perspective API set $S^{*}(f_{C})$ satisfied JR 30%, 50%, and 100% of the time, respectively. This was surprising as the Perspective API does not explicitly aim to satisfy JR. Moreover, both diverse approval and the Perspective API classifiers are meant to“bridge” across groups (Ovadya & Thorburn, 2023). However, Perspective API satisfied JR at a much higher rate and also provided better overall representation (Fig. 2). This could be because diverse approval focuses on bridging across a few pre-defined groups (in this case, the three political groups), whereas the bridging attributes targeted by the Perspective API might appeal to a broader range of groups. [^15] All our results rely on Remesh’s inferred approval votes and biases in these inference could impact our results. For example, if the inferences are biased towards predicting approval for longer texts (we do not know if this is the case), which are also scored higher by the Perspective API, then the Perspective API might appear to be provide more representation than it actually does.
@@
[…]
The JR feeds also satisfied EJR+. Finally, out of the 48 JR feeds in our experiment, 47 of them also satisfied EJR+ (Brill & Peters, 2023), a strengthened version of JR. Although more experimentation is needed, these findings may indicate that for comment ranking, as in other real-world applications of JR (Boehmer et al., 2024), the practical differences between these axioms may be minimal.
@@ 7 Discussion
Limitations. Nevertheless, there are still additional questions that should be considered before deploying our framework on a real-world platform.
First, the JR representation constraint depends on an approval matrix specifying which users approve which comments. It is not always clear which users should be included in the approval matrix. For example, should it be all users on the platform or only those who saw the post? The answer to this question may be different for different platforms.
@@
Future work. In this work, we focus on analyzing JR (although empirically, we find that all but one of our JR committees also satisfies EJR+). In future work, it would be interesting to analyze stronger axioms such as EJR (Aziz et al., 2017), EJR+ (Brill & Peters, 2023) or BJR (Fish et al., 2024). Additionally, our work focused on the classical JR setting, where a set of $k$ top comments is selected. However, it could be interesting to explore extensions of JR that apply directly to the ranking setting (Skowron et al., 2017; Israel & Brill, 2024), although it is not immediately clear how to integrate a scoring function into these extensions.
@@ Acknowledgements
We thank Hoda Heidari, Seth Lazar, Aviv Ovadya, Ariel Proccacia, Luke Thorburn, and Glen Weyl for their feedback on this project. We thank Andrew Konya and Lina Qiu for sharing the Remesh inferred approval votes with us and making them publicly available for future research.
Impact Statement
Online comment sections have the potential to serve as important platforms for public discourse and deliberation, though their effectiveness is often undermined by the low quality of conversations. Algorithms can potentially significantly improve these spaces by facilitating more constructive public sphere deliberation. This work takes a step in this direction by directly incorporating ideals of representation, as formalized in social choice theory, into the comment ranking process. We demonstrate that this approach can maintain conversational quality while ensuring diverse viewpoints are represented. Nonetheless, as discussed in our limitations section, before deploying this on a real-world platform, more testing is necessary to ensure that implementations of this approach genuinely provide users with representation.
@@ Appendix A How does JR compare to demographic-based representation?
It is worth discussing how our JR-based approach differs from other, perhaps more familiar, notions of demographic or social representation from the algorithmic fairness literature (Barocas et al., 2023; Chasalow & Levy, 2021). JR is a bottom-up measure where the groups that are represented depend entirely on the users’ own approval of different comments. This has several advantages for the comment ranking setting.
First, is the feasibility for real-world applications. In algorithmic fairness, groups are typically pre-defined based on a demographic, such as race or religion. In contrast to JR, the requirement of (inference of) user demographic labels make applying many algorithmic fairness methods infeasible or difficult to implement in industry settings, including on social media platforms, because of conflicts with privacy and legal constraints (Holstein et al., 2019; Veale & Binns, 2017).
@@
Third, JR may accommodate intersectionality (Crenshaw, 1991) in a more natural way than many algorithmic fairness approaches. Algorithmic fairness approaches to intersectionality attempt to provide guarantees to a wide set of subgroups (Gohar & Cheng, 2023), e.g., by defining intersectional groups as the combination of different demographic attributes (e.g. age, gender, race) (Kearns et al., 2018). However, this approach can lead to issues where the subgroups no longer correspond to meaningful entities. For instance, intersecting many dimensions can result in subgroups that are too specific and lack a meaningful reason for grouping (e.g., ”Jewish white males aged 18-35 who have a bachelor’s degree and live in a rural area”). In contrast, JR focuses on cohesive groups—groups of users who can agree on at least one comment in common—thereby inherently embedding some requirement for meaningfulness.
@@ Appendix B An Example of Conflict between Diverse Approval and JR
Fig. B.1: When diverse approval fails at representativeness An approval profile with $n=12,m=5$ and $k=3.$ There are $\gamma=2$ cohesive groups $G_{1}=\{A,B,C,D,E,F\}$ and $G_{2}=\{G,H,I,J,K,L\}$ such that $|G_{1;2}|=6>n/k.$ Note that the sub-groups such as $\{A,B,C,D\}$ and $\{I,J,K,L\}$ are also cohesive and larger than $n/k$, so they ought to also be represented by JR. In turn, any JR set $S$ contains $\{1,2\}.$ On the other hand, the comments’ maximin diverse approval scores (across the groups $G_{1}$ and $G_{2}$) are $f(1)=f(2)=0$ and $f(3)=f(4)=f(5)=1/6.$ In turn, the set with the highest score is $S^{*}=\{3,4,5\}.$
@@ C.2 Proof of Theorem 1
We first prove the following lemma. This statement can be found in (Bredereck et al., 2019) (Theorem 3), but the authors do not provide a proof for the result in their paper so we write one below.
@@ C.4 Proof of Theorem 3
There are several generative interpretations of the Mallows model (Lu & Boutilier, 2014) but one that is particularly useful for analyzing homogeneity is based on the Repeated Insertion Model (RIM) (Doignon et al., 2004). In our analysis, we make use of a modified version of RIM in which items from the reference ranking are inserted starting with the last ranked item and ending with the first ranked item. It can be shown this bottom up RIM sampling induces the same distribution as the usual top down RIM sampling.
@@ D.1 The GreedyCC approximation algorithm
First, we review the GreedyCC algorithm used in both sets of experiments. Given that solving for the optimal JR set $S^{*}_{JR}$ is NP-hard in general (Bredereck et al., 2019; Elkind et al., 2022), we would expect practitioners to resort to approximation algorithms. Thus, to understand the effects we might expect in practice, in all our experiments, we too employ an approximation algorithm, along with an associated approximate price of JR. Specifically use the GreedyCC algorithm, which has also been used in prior work (Elkind et al., 2022), to find a set $S^{\text{Greedy}}_{JR}$ that satisfies JR and approximately maximizes the score function $f$. In brief, our implementation of GreedyCC greedily adds items to the selected set until a $n/k$-justifying set is found; the remaining items are then selected to maximize the score function $f$. A full description of the algorithm is outlined in Alg. 18.
@@ D.2 Mallows Mixture Model Simulations
$\mathcal{M}_{2}([\phi,\phi],[\sigma_{1},\sigma_{2}],[1/2,1/2])$ that simulates a polarized environment with two groups. We draw inspiration from Esteban & Ray (1994)’s axiomatic framework for measuring polarization. Esteban & Ray (1994) where they consider polarization to be present if three intuitive properties are satisfied. We operationalized these three properties in our simulation as follows:
@@ Results.
Second, we find that the price of GreedyCC is higher for diverse approval than for engagement. For diverse approval, some instances result in a price of GreedyCC that nearly our theoretical bound from Theorem 3. The gap in price between diverse approval and engagement is consistent with our result in Theorem 1, which found that the price of JR for maximin diverse approval $P(k,f_{\text{MDA}})$ equals $k$, whereas Elkind et al. (2022) demonstrated that the price for engagement $P(k,f_{\text{eng}})$ is $\Theta(\sqrt{k})$. These results were shown, in the general setting, without any restrictions on the approval profiles. On the other hand, the theoretical bound we derived for the polarized setting in Theorem 3, and plotted in Figure D.1, holds for all score functions. But given the empirical disparity in price between diverse approval and engagement, it seems plausible that a similar score-function-specific gap could be identified even in the polarized setting.
@@ E.2 Perspective API Classifiers
Fig. E.1: Scores from the Perspective API
@@ E.4 Size of the $n/k$-justifying Set Found by GreedyCC
In Section 5.2, we show that when the user population clusters into a few groups, then the price of JR will be low because it is possible to find a small $n/k$-justifying set (and thus, the remaining items for the set can be chosen to maximize the score function). Indeed, this logic is how the GreedyCC approximation algorithm that we use operates. As detailed in Footnote 18, the GreedyCC proceeds in two stages. In the first stage, it finds an $n/k$-justifying set. In the second stage, it selects the remaining comments that maximize the score function. Thus, the smaller the $n/k$-justifying set, the lower the price of GreedyCC will tend to be.